OpenStudy (summersnow8):

Five points, A,B,C,D,E, are equally spaced, in that order, on the circumference of a circle wither center O. Let X be the intersection of chords AC and BD. Find the 3 angles of triangle ABX. [Hint: the measure of an angle inscribed in a circle (for example, angle BAC) is 1/2 the measure of the central angle that subtends the same arc (angle BOC)

OpenStudy (anonymous):

 We know that the angle between any two points through the origin is 360/(number of points)=360/5=72. From that we can solve the rest.

OpenStudy (summersnow8):

I don't really understand, but thanks :/